I really need some help with this. Please. Find the axis of symmetry.
Y = x^2 + 5x -9
Thanks

Answer by stanbon(48558)

Find the axis of symmetry.
Y = x^2 + 5x -9
---------
There are at least two ways to do this:
1st: Find the x-value of the max point at x=-b/2a
x=-5/2
That is the axis of symmetry.
-------
2nd: Put in vertex form:
y+9 = x^2+5x
Complete the square to get:
x^2+5x+(5/2)^2 = y+9+(5/2)^2
Factor to get:
(x+(5/2))^2 = y+(61/4)
This tells you the vertex is at ((-5/2),(-61,4))
Axis of symmetry is x=-5/2
========
Cheers,
Stan H.